OpenStudy (anonymous):

(1,19),(2,4); y=a(x+1)^2+k determine a and k so the given points are on the graph of the function

OpenStudy (anonymous):

plug in two point in x, y to find a and k

OpenStudy (anonymous):

19=a(1+1)^2+k 4=a(2+1)^2+k

OpenStudy (anonymous):

replace x by 1 set result equal 19, get an equation in a and k replace x by 2 set result equal 4, get another equation in a and k. solve the system of equations

OpenStudy (anonymous):

what moneybird said

OpenStudy (anonymous):

\[4a+k=19\] \[9a+k=4\] \[-5a=15\] \[a=-3\] etc

OpenStudy (anonymous):

19 = 4a + k 4 = 9a + k 4 = 9a + k - 19 = 4a + k ------------ -15 = 5a a = -3 k = 31